Calculate E for the following electrochemical cell at 25 oC given the following standard reduction potentials. Agl(s) e- Ag(s) I(aq) Eo -0.15 V Sn4 (aq) 2 er Sn2 (aq) Eo +0.15 V
Answer #1

Sn is anode and AgI is cathode.

So, oxidation occurs at anode and reduction at cathode.

Eo = Eanode + Ecathode

= (-0.15) + (-0.15) = -0.30 V

EMF = Eo - 0.06/n log [I-]2[Sn4+]/[Sn2+]

= -0.30 - 0.06/2 log (0.15)2(0.50)/(0.50)

= -0.30 - 0.03 log (0.15)2

= -0.30 - 0.03 log (0.0225)

= -0.30 - (0.03)(-1.65)

= -0.30 + 0.05 = -0.25 V

E of the cell is -0.25 V